OpenStudy (anonymous):

According to the general equation for conditional probability, if p (a^b')=1/6 and p(b')=7/24 , what is p(a[b) ? 3/7 4/7 5/7 2/7

OpenStudy (anonymous):

@pooja195 @Ms-Brains

OpenStudy (anonymous):

@Ms-Brains hello

MsBrains (ms-brains):

Hey, First we need to convert this in to a number problem. P(a^b) =1/6 and P(b) = 7/24 What is p(a[b) Figure this out first.

OpenStudy (anonymous):

how would we convert it

MsBrains (ms-brains):

I am not sure, I mean, I never learned this. I actually have no idea if that is what we are supposed to do. ;-; Sorry buddy.

OpenStudy (anonymous):

lmao wut xDD

OpenStudy (anonymous):

it looked like u knew what u were doin in ur first response

OpenStudy (anonymous):

lmao u funny xDDD

MsBrains (ms-brains):

Yup, :P I'm sorry, I thought that was how you solve it. Maybe somebody who knows this can help. cx